# A Hierarchical Bin Packing Framework with Dual Manipulators via Heuristic Search and Deep Reinforcement Learning

### v1

@@ V-B Training Results
V-C Algorithm Tests
V-C1 Scenarios
TABLE II: Packing performance across different scenarios.
@@ V-C1 Scenarios
(a) Performance on Random instances.
@@
(b) Performance on 100% set with and without repacking.
We define six distinct scenarios to evaluate our bin packing framework under varying robotic configurations and perception constraints, as summarized in Table I and shown in Fig. 1. Each scenario label encodes the following:
@@
  • A: number of accessible items per robot,
@@
  • O: number of items in the overlapping workspace (for dual-robot settings only).
@@
Scenarios S-R1A1, S-R5A1, and S-R5A3 represent single-arm configurations, while D-R2A2O2, D-R5A2O2, and D-R5A3O1 correspond to dual-arm settings. S-R1A1 denotes a minimal setup where one robot can recognize and access only a single item. In contrast, D-R5A3O1 reflects the most capable configuration, where two robots can recognize five items, each access three, and share one item in the overlapping region.
V-C2 Test with Random Instances
@@
We evaluate our method on the Random instances with 1000 episodes per scenario, where each episode involves packing a single bin. All scenarios share the same item arrival sequence across the entire test. We first evaluate our method without repacking across six scenarios, with and without item rotation. As shown in Table II(a), both bin utilization and the total number of packed items consistently improve when rotation is enabled. For instance, in Scenario S-R1A1, the utilization increases from 78.88% to 85.17%, and the number of packed items from 8715 to 9422. Similar trends are observed across all scenarios, with Scenario D-R5A3O1 achieving the highest utilization of 98.59%. We then test repacking in four scenarios where there is room for improvement: S-R1A1, S-R5A1, D-R2A2O2, and D-R5A2O2. Fig. 8 shows that even a short planning time of 1 second for repacking significantly improves the final bin utilization. In Scenario S-R1A1, the utilization increases from 85.17% to 98.01%, and in D-R2A2O2 from 91.57% to 99.03%. In Scenario S-R5A1, the utilization improves from 92.38% to 99.20%, and in D-R5A2O2 from 96.46% to 99.67%. Scenarios with more known items result in fewer repacking actions, which helps reduce motion overhead during execution.
@@
V-C3 Test with 100% set
We evaluate our method on the 100% set with 3000 episodes per scenario, where each episode uses the same fixed item arrival sequence. As shown in Table II(b), our method achieves high utilization even without repacking. When repacking is enabled, 100% utilization is achieved in all scenarios. The number of fully packed bins increases significantly with repacking, from 684 to 3000 in S-R1A1 and from 865 to 3000 in D-R2A2O2. Repacking improves packing quality at the cost of moderate computation time, with an average of less than 3 seconds to find a repacking solution.
V-C4 Compare with Baseline
TABLE III: Comparison of bin utilization (%) with baseline methods in $4\times 4$ and $5\times 5$ bin environments.
@@ V-C4 Compare with Baseline
To validate the effectiveness of our method, we compare it with existing baselines, including heuristic methods (Skyline [4], ShelfNextFit [5]) and learning-based approaches (Deep-Pack [7], BERM [19]). The evaluation follows the same setup as the baselines in $4\times 4$ and $5\times 5$ bin environments under a fully online setting without rotation or repacking, ensuring fair comparison. Table III shows that our method achieves the highest mean bin utilization in both settings—98.23% in the $4\times 4$ bin (compared to 94.87% for Deep-Pack and 97.23% for BERM), and 96.58% in the $5\times 5$ bin (compared to 91.48% and 94.30%, respectively). These results demonstrate that our DRL-based policy outperforms the state-of-the-art methods in a constrained online scenario.
V-D Simulation Tests
The simulation environment is developed using Unity [28] to evaluate the proposed framework under physical constraints. Universal Robots UR5e is used as the manipulator in both single- and dual-arm configurations. The simulation environments for each setting are shown in Fig. 9. For motion planning, cuRobo [29] is employed in both configurations, while StopNGo [13] is integrated to coordinate dual-arm execution. Each setup is tested over 30 episodes using the 100% set instances.
@@ V-D Simulation Tests
As shown in Table IV, the dual-arm scenario achieves a shorter makespan than the single-arm scenario, despite requiring longer per-step planning durations. Specifically, it reduces the overall makespan by approximately 24% in both repacking and non-repacking cases. Since planning and execution proceed concurrently, planning for the next step is typically completed during the current execution, minimizing idle time. Although repacking introduces slight overhead in planning and execution, it improves bin utilization (as discussed in Sec. V-C3), which justifies the additional cost. In instances with an odd number of items, the final item is always handled by a single robot, limiting the relative gain in makespan reduction. These results demonstrate that the proposed framework can generate physically feasible and time-efficient plans for dual-arm robotic systems.
(a) S-R5A3 simulation environment
(b) D-R5A3O1 simulation environment
Figure 9: Simulation environments used for testing. (a) single-arm setting (S-R5A3), and (b) dual-arm setting (D-R5A3O1).
TABLE IV: Planning and Execution Time in Single- and Dual-Arms
@@ VI Conclusion
This paper presented a hierarchical bin packing framework that combines heuristic tree search and deep reinforcement learning to solve 2D bin packing problems. The method supports repacking and dual-arm coordination through integrated task and motion planning. A low-level RL agent proposes placement positions, while the high-level search explores packing and unpacking strategies under item availability and accessibility constraints. Experiments demonstrate near-optimal bin utilization and show that the framework generates physically feasible, time-efficient plans, reducing execution time in dual-arm scenarios. These results highlight the benefits of repacking and parallel execution for practical robotic systems.

### v2

@@ V-B Training Results
TABLE III: Packing performance across different scenarios.
@@
(a) Performance on Random instances
@@
(b) Performance on 100% set with and without repacking
V-C Test Environment
V-C1 Scenarios
We define six distinct scenarios to evaluate the proposed bin packing framework under varying robotic configurations and perception constraints, as summarized in Table II and illustrated in Fig. 1. Each scenario label is encoded as follows:
@@ V-C1 Scenarios
  • A: maximum number of accessible items per robot,
@@
  • O: maximum number of items in the overlapping workspace (for dual-robot settings only).
@@
Scenarios S-R1A1, S-R5A1, and S-R5A3 represent single-robot configurations, while D-R2A2O2, D-R5A2O2, and D-R5A3O1 correspond to dual-robot settings. S-R1A1 denotes the minimal case, where a single robot can recognize and access only one item at a time. In contrast, D-R5A3O1 represents the most capable configuration, where two robots can recognize five items, each can access three items, and one item lies within the overlapping workspace shared by both robots. The maximum number of items that can be accessed by all robots, denoted as $n_{\max}$, is given by $n_{\max}=n_{\text{A}}\times n_{\text{robot}}-n_{\text{O}}\times(n_{\text{robot}}-1)$, where $n_{\text{A}}$ denotes the number of accessible items per robot, $n_{\text{O}}$ represents the number of items located in the overlapping workspace, and $n_{\max}\leq n_{\text{R}}$, with $n_{\text{R}}$ being the number of recognized items.
V-C2 Conveyor Logic
Since our search algorithm considers both robot accessibility and task allocation, the conveyor belt logic for positioning incoming items is a critical component of the experimental setup. For single-robot scenarios or dual-robot configurations where both robots share the same accessible region (e.g., $n_{\text{A}}=n_{\text{O}}$ in D-R2A2O2 and D-R5A2O2), items can be continuously positioned along the conveyor. However, in configurations such as D-R5A3O1, where the maximum number of items accessible to each robot ($n_{\text{A}}$) differs from that in the overlapping region ($n_{\text{O}}$), exclusive zones arise that can be reached by only one robot. If more items accumulate in one exclusive zone, the advantages of using dual robots diminish due to workload imbalance. To maintain balanced item allocation, we model the conveyor control logic at each time step as follows.
Notation:
  • $k_{\text{total}}$: Total number of items on the conveyor
  • $k_{\alpha}$: Number of items in the exclusive region of the rear robot
  • $k_{\beta}$: Number of items in the exclusive region of the front robot
  • $k_{\text{O}}$: Number of items in the overlapping region (accessible to both robots)
  • $t$: Time step before conveyor operation
  • $t+1$: Time step after conveyor operation
When $k_{\text{total},t+1}\geq n_{\text{max}}$, where $n_{\text{max}}=2n_{\text{A}}-n_{\text{O}}$ denotes the maximum number of items that both robots can collectively access, items are continuously positioned along the conveyor as in the fully shared-region cases. When $k_{\text{total},t+1}<n_{\text{max}}$, the item distribution across regions is determined as follows:
$k_{\text{O},t+1}$ $=\min\!\left(k_{\text{total},{t+1}}-k_{\alpha,t},\,n_{\text{O}}\right),$ (7)
$k_{\alpha,{t+1}}$ $=\max\!\left(k_{\alpha,t},\,\left\lfloor\frac{k_{\text{total},t+1}}{2}\right\rfloor+1-k_{\text{O},t+1}\right),$ (8)
$k_{\beta,{t+1}}$ $=k_{\text{total},t+1}-k_{\alpha,t+1}-k_{\text{O},t+1}.$ (9)
By controlling the conveyor belt to place items sequentially into the $\alpha$, O, and $\beta$ regions according to the computed quantities, the system maintains a balanced allocation of items between the workspaces of the two robots even when their accessible regions differ, thereby preserving the efficiency gains of dual-robot operation. All following tests are conducted under the assumption that the conveyor belt operates according to the described logic at every time step.
V-D Algorithm Tests
Before discussing each result in detail, we briefly summarize the purpose and scope of the ablation studies conducted to verify the contribution of each component in the proposed hierarchical framework. In summary, the ablation results demonstrate that:
  • under the consideration of rotation and multiple perceived items, the tree search component improves orientation selection and packing efficiency;
  • the dual-manipulator setting significantly reduces packing steps while maintaining high utilization;
  • the tree search handles multiple reachable items and performs task allocation, especially in dual-manipulator settings;
  • the repacking algorithm further enhances bin utilization across all scenarios within a limited planning time;
  • the reinforcement learning module alone achieves higher utilization than the baseline algorithms.
Detailed analyses of each case are presented in the following subsections.
V-D1 Test with Random Instances
We evaluate our method on the Random instances with 1000 episodes per scenario, where each episode involves packing a single bin. In the Random instance, the items that has not been packed at the current episode, are not discarded but are reconsidered in subsequent planning stages. Such items are revisited either during the repacking phase or when planning for the next bin, ensuring that all items have the opportunity to be packed. This cumulative process over 1,000 bins is reflected in the total number of packed items reported in Table III(a). All scenarios share the same item arrival sequence across the entire test.
We first evaluate our method without repacking across six scenarios, with and without item rotation. When performing packing only, the Selection described in Alg. 2 (line 18) is skipped, and the total planning time for all scenarios averages less than 1.2 seconds. As shown in Table III(a), both bin utilization and the total number of packed items consistently improve when rotation is enabled. For instance, in Scenario S-R1A1, the utilization increases from 78.88% to 85.17%, and the number of packed items from 8715 to 9422, achieving higher throughput within the same number of bins. Similar trends are observed across all scenarios, with Scenario D-R5A3O1 achieving the highest utilization of 98.59%. Furthermore, the influence of the number of perceived items on performance is verified by comparing S-R1A1 vs. S-R5A1 and D-R2A2O2 vs. D-R5A2O2. In addition, when comparing S-R5A1 vs. S-R5A3 and D-R5A2O2 vs. D-R5A3O1, where each robot has more reachable objects, the results show consistent improvements in utilization. In the dual-arm cases (D-R5A2O2 vs. D-R5A3O1), even with a larger number of packed items, the total number of packing steps decreased. These results collectively verify that the proposed candidate-selection strategy is valid and that the method successfully performs task allocation between the two manipulators, directly indicating improved throughput per cell.
@@ V-D1 Test with Random Instances
We then test repacking in four scenarios where there is room for improvement: S-R1A1, S-R5A1, D-R2A2O2, and D-R5A2O2. In these repacking experiments, the number of candidates selected by the Selection function is fixed to two to compare the utilization achieved under a limited additional planning time for repacking. Fig. 8 shows that even a short planning time of 1 second for repacking significantly improves the final bin utilization. In Scenario S-R1A1, the utilization increases from 85.17% to 98.01%, and in D-R2A2O2 from 91.57% to 99.03%. In Scenario S-R5A1, the utilization improves from 92.38% to 99.20%, and in D-R5A2O2 from 96.46% to 99.67%. The fact that such high improvements are achieved within a short planning time indicates that the proposed last-packed-first search strategy efficiently identifies promising repacking policies. Scenarios with more known items result in fewer repacking actions, which helps reduce motion overhead during execution.
@@
Figure 9: Repacking success count and planning time with varying numbers of candidates selected by the Selection in D-R5A3O1 on the 100% set.
V-D2 Test with 100% set
We evaluate our method on the 100% set with 3000 episodes, where each episode uses the same fixed item arrival sequence per scenario. In the 100% set experiments, items that have no valid placement are re-evaluated during the repacking phase, and if infeasible placements remain, those items are excluded from further consideration, since the item sets are predefined for each bin.
When performing packing only, the Selection in Alg. 2 (line 18) is skipped, and the total planning time for all scenarios averages less than 1.3 seconds. As shown in Table III(b), our method achieves high utilization even without repacking. Furthermore, performance improves as the number of perceived and reachable items increases, confirming that the hierarchical decision process leverages additional information.
In addition, we conduct an experiment on the 100% set to measure the repacking success count and computation time with varying numbers of candidates selected by the Selection function in D-R5A3O1, as presented in Fig. 9. The Selection process is applied until the number of remaining items in the search tree drops below three; when fewer than three items remains, pruning is skipped to preserve remaining candidates, since the search space becomes sufficiently small due to additional early-termination conditions implemented in Alg.2 (lines 10–11,29–30). When repacking is enabled with this setting, 100% utilization is achieved across all scenarios. The number of fully packed bins increases significantly with repacking, from 684 to 3000 in S-R1A1 and from 865 to 3000 in D-R2A2O2. Repacking improves packing quality at the cost of moderate computation time, with an average of less than 3 seconds to find a repacking solution. These findings further confirm that the proposed hierarchical framework achieves near-optimal utilization under realistic time constraints.
V-D3 Compare with Baseline
TABLE IV: Comparison of bin utilization (%) with baseline methods in $4\times 4$ and $5\times 5$ bin environments.
@@ V-D3 Compare with Baseline
To validate the effectiveness of our method, we compare it with existing baselines, including heuristic methods (Skyline [4], ShelfNextFit [5]) and learning-based approaches (Deep-Pack [7], BERM [26]). The evaluation follows the same setup as the baselines in $4\times 4$ and $5\times 5$ bin environments under a fully online setting without rotation or repacking, ensuring fair comparison. Table IV shows that our method achieves the highest mean bin utilization in both settings—98.23% in the $4\times 4$ bin (compared to 94.87% for Deep-Pack and 97.23% for BERM), and 96.58% in the $5\times 5$ bin (compared to 91.48% and 94.30%, respectively). Assuming that each pallet or tote corresponds to a fixed container volume, this improvement directly translates to a reduction of roughly one pallet for every 20–40 bins packed, resulting in measurable logistics cost savings and reduced storage requirements. The superior performance of our method can be attributed to the DRL formulation. Compared with heuristic-based baselines [5, 4], which usually follow sequential placement heuristics, our RL agent is free to explore and select placements from any direction within the bin, allowing more flexible decision-making. Furthermore, our method surpasses existing DRL-based algorithms [7, 26] due to two main design choices:
  • the use of diverse training datasets that improve generalization across different item distributions,
  • a reward formulation defined as the number of occupied cells adjacent to the packed item (described in Sec. IV-A), which provides dense and flexible rewards and induces generalizable packing behaviors.
V-E Simulation Tests
V-E1 Configuration of Simulation Environment
The simulation environment is developed using Unity [30] to evaluate the proposed framework under physical constraints. Universal Robots UR5e is used as the manipulator in both single- and dual-arm configurations. A square bin of 45 cm $\times$ 45 cm is used, discretized into a $10~\times~10$ grid where each cell corresponds to 4.5 cm. The simulation environment layout for the dual-robot scenario, and the packing scenes of single-arm and dual-arm configurations are illustrated in Fig. 11. The design of the simulation environment is primarily guided by the workspaces of both manipulators. Each robot is positioned such that both arms can fully cover the bin area while maintaining sufficient reach to grasp incoming items on the conveyor belt. To prevent mutual motion interference, the robots are spaced approximately half of their reach distance apart, which reflects safety and accessibility considerations. For the single-robot configuration, only the right arm is used to ensure consistent comparison under identical environmental conditions.
Fig. 10 illustrates the action sequence of the robot during the packing process. The interval between incoming items on the conveyor belt is set to 0.7 m, and the conveyor speed is configured to 0.5 m/s. When the conveyor operates and a new item arrives, the planning module generates a placement plan, after which the robot initiates the packing task. The packing operation consists of repeated pick-and-place actions. Once the picking of the current item on the conveyor is completed, the conveyor operates again to transport the next item, thereby allowing conveyor operation and robotic placement to proceed in parallel. Once the new item has arrived and the planning for the next step is completed, followed by the completion of the previous placement, the next operation begins (if repacking is required, the repacking phase is executed first).
V-E2 Makespan Evaluation
In this paper, we define the makespan as the total time required for the robot(s) to complete the loading process within a single episode, including conveyor operation, planning, and execution, as shown in Fig. 10. For motion planning, cuRobo [31] is employed in both configurations, while Stop-N-Go [13] is integrated to coordinate dual-arm execution. Each setup is tested over 30 episodes using the 100% set instances. The average time between the conveyor activation and the arrival of a new item is approximately 1.7 seconds.
@@ V-E2 Makespan Evaluation
As shown in Table V, the dual-arm scenario achieves a shorter makespan than the single-arm scenario, despite requiring longer per-step planning durations. Specifically, it reduces the overall makespan by approximately 24% in both repacking and non-repacking cases, corresponding to a proportional increase in cell throughput under identical task conditions. Since planning and execution proceed concurrently, planning for the next step is typically completed during the current execution, thereby minimizing idle time. Although repacking introduces slight overhead in planning and execution, it improves bin utilization (as discussed in Sec. V-D2), which justifies the additional cost. In instances with an odd number of items, the final item is always handled by a single robot, limiting the relative gain in makespan reduction. These results demonstrate that the proposed framework can generate physically feasible and time-efficient plans for dual-arm robotic systems.
V-F Physical Robot Test
In the physical robot experiments, a 32 cm $\times$ 26 cm bin is used with 1 cm margins on both sides, corresponding to a 3 cm grid resolution. Two AgileX Piper manipulators [32] are employed for dual-arm operation, following the D-R5A3O1 scenario setting. ArUco markers are attached to the objects for vision-based detection and pose estimation. The experimental setup and dual-arm operation are illustrated in Fig. 12. Motion planning for dual manipulators is performed using cuRobo [31] and Stop-N-Go [13]. Through these experiments, we demonstrated the feasibility of applying the proposed framework to real dual-arm packing tasks in a logistics environment.
Figure 10: Action sequence and makespan of single-arm and dual-arm configurations.
(a) Layout of the dual-arm simulation environment (D-R5A3O1)
(b) Single-arm packing scene (S-R5A3)
(c) Dual-arm packing scene (D-R5A3O1)
Figure 11: Overview of the simulation environment and packing scenes. In (a), the layout of the dual-arm simulation environment is shown, where the circles indicate the workspace of each manipulator, while (b) and (c) show the packing scenes of the single-arm and dual-arm configurations, respectively.
TABLE V: Planning and Execution Time in Single- and Dual-Arms.
@@ V-F Physical Robot Test
Figure 12: Real dual-arm packing experiment setup showing the manipulators, bin, and vision-based detection system.
@@ VI Conclusion
This paper presented a hierarchical bin packing framework that combines heuristic tree search and deep reinforcement learning to solve 2D bin packing problems. The method supports repacking and dual-arm coordination through integrated task and motion planning. A low-level RL agent proposes placement positions, while the high-level search explores packing and unpacking strategies under item availability and accessibility constraints. Experiments demonstrate near-optimal bin utilization and show that the framework generates physically feasible, time-efficient plans, reducing execution time in dual-arm scenarios. These results highlight the benefits of repacking and parallel execution for practical robotic systems. Future work is left for extending the current 2D framework to 3D by replacing the binary grid matrix representation with a height-map formulation, while incorporating dual-arm workspace constraints for three-dimensional packing tasks.